Find the value(s) of k for which the equation (x+2)(x+k)=-1 has equal roots.
Answer by stanbon(75887)

Find the value(s) of k for which the equation (x+2)(x+k)=-1 has equal roots.
-------
x^2+(k+2)x+2k = -1
---
Has equal roots when the discriminant equals zero.
---
(k+2)^2 -4*(2k) = 0
---
k^2+4k+4-8k = 0
---
k^2-4k+4 = 0
(k-2)^2 = 0
k = 2
================
Cheers,
Stan H.